BINDING ENERGIES OF Po AND Po^{+} WITH ATMOSPHERIC NITROGEN

\underline{T. M. MOFFETT} and P. A. CHRISTIANSEN, Department of Chemistry, Clarkson University, Potsdam, NY 13699-5810.

The interactions between polonium, a decay product of radon, and nitrogen were examined in an ab-initio study. Properties (R_{e}, D_{e}) of linear and cyclic N_{2}Po and N_{2}Po^{+} were determined at various levels of theory (HF, MP2, MRCI including spin-orbit coupling). The geometric arrangements with the lowest energies were found to be linear for both the neutral molecule and the positive ion. Binding energies in both species were determined to be approximately 0.1 eV.